Ritz​ Products's materials​ manager, Tej​ Dhakar, must determine whether to make or buy a new semiconductor for the wrist TV that the firm is about to produce.

Two million units are expected to be produced over the life cycle. If the product is​ made, start-up and production costs of the make decision total

​$1 million, with a probability of 0.4 that the product will be satisfactory and a 0.6 probability that it will not. If the product is not​ satisfactory, the firm will have to reevaluate the decision. If the decision is​ reevaluated, the choice will be whether to spend another​ $11 million to redesign the semiconductor or to purchase. Likelihood of success the second time that the make decision is made is 0.7

If the second make decision also​ fails, the firm must purchase. Regardless of when the purchase takes​ place, Dhakar's best judgment of cost is that Ritz will pay​$0.50

for each purchased semiconductor plus​$1 million in vendor development cost.​a) Assuming that Ritz must have the semiconductor​ (stopping or doing without is not a viable​ option), what is the best​ decision?

The firm should (make or buy?) the semiconductors because this decision has an expected cost of ________

What criteria did you use to make this​ decision?

Note that we determined the expected monetary value for each decision in part​ (a).

In this​ case, expected monetary value is represented by _______?

What is the worst that can happen to Ritz as a result of this particular​ decision? What is the best that can​ happen?
Answer #1

Buying cost = (2x0.5+1 )M = 2M

making cost =

Cost of make option = 1M+ 0.6 [ 11M+0.3x2M} = 7.96M

(a) Firm should buy because the cost of buying is 2M against 7.96M of making option.

(b) EMV of successful decision is a cost of 2M.

(c) The best which can happen to the company is success in the first trial, which has a cost of 1M.

The worst which can happpen to the company is that the company fails in two attempts and then buys the item

total cost of this option = 1M+11M+2M = 14M